. Calculating tax incidence suppose that the u·s·government decides to charge oola consumers a tax. Before the tax, 50,000 cases ofcos were sold every week at a price of $7 per case. After the tax, 44,000 cases of cola are sold every week; consumers pay $10 per case (inluding the tax), and producers receive $4 per case. of this amount, the burden thast tails on consumers is per case, and the The amount of the tax on a case of cola is burden that fals an praducers s S per casc True or False: The effect of the tax on the quantity sold would have been the same as the tax had been levied on producers. O True False
Answer #1

Amount of tax=consumer pay-seller received=10-4=$6

consumer burden=$10-7=$3

and producer burden =$7-$4=$3

true because end result doesnot depend on which party the tax is imposed